A student knows 20 answers from the 25 given questions of the chemistry exam.
What is probability that student will answer not less than 2 questions from 3 ones
taken at random?
1
Expert's answer
2020-11-08T18:56:07-0500

The probability that student will answer not less than 2 questions from 3 ones

taken at random is:

"P(2, 3) = \frac{3!} {2!×(3 - 2)!}×0.8^2×0.2^1 + 0.8^3 = 0.384 + 0.512 = 0.896."